How much work does the electric field do in moving a proton from a point with a potential of +130 V to a point where it is -70 V ? Express your answer in joules.

Express your answer in electron volts.
Answer #1

Workdone by electric field, W = F*d

= q*E*d

= q*delta_V

= q*(V1 - V2)

= 1.6*10^-19*(130 - (-70))

= 3.2*10^-17 J